# DPZV: Elevating the Tradeoff between Privacy and Utility in Zeroth-Order Vertical Federated Learning

### v2

@@ 3 Background and Motivation
One of the most widely adopted private training algorithms is Differentially Private Stochastic Gradient Descent (DP-SGD), which operates by clipping an intermediate result and injecting Gaussian noise during each model update. Specifically, given a minibatch $B_{t}$ at iteration $t$, DP-SGD performs the following steps: 1) Clip individual gradients: $\bar{g}_{i}=\nabla\ell(\theta_{t},x_{i}),\quad\tilde{g}_{i}=\bar{g}_{i}/\max\left(1,\frac{\|\bar{g}_{i}\|_{2}}{C}\right)$. 2) Add Gaussian noise: $\tilde{g}_{t}=\frac{1}{|B_{t}|}\left(\sum_{i\in B_{t}}\tilde{g}_{i}+\mathcal{N}(0,\sigma^{2}C^{2}\mathbf{I})\right)$. 3) Update model: $\theta_{t+1}=\theta_{t}-\eta_{t}\tilde{g}_{t}$.
@@
Challenges and Motivation. While DP-SGD is effective in centralized or horizontally federated setups, it becomes more challenging in VFL, particularly when ZO optimization is employed. First, due to the distributed nature of VFL, differential privacy must be enforced before the server communicates with clients to protect against potential inference attacks by malicious client parties. In both cases, the noise must match the dimensionality of the protected vector (embedding or gradient), often resulting in high-dimensional noise. This situation is further exacerbated by the second factor: in ZO methods, gradients are approximated through multiple forward evaluations, making them inherently noisy. Adding a high-dimensional noise vector as in standard DP-SGD further amplifies this inaccuracy.
@@
A notable benifit of ZO optimization is its communication efficiency in the backward pass, where only a scalar value instead of high-dimensional gradients are transmitted. This enables a more efficient privacy mechanism—injecting scalar noise rather than high-dimensional noise. Motivated by this observation, we develop Differentially Private Zeroth-Order Vertical Federated Learning (DPZV), a framework that achieves tunable differential privacy through calibrated scalar noise injection. This design allows for a favorable privacy-utility trade-off in ZO-based VFL, particularly under tight privacy budgets.
4 Methodology
@@
In this section, we describe the overall training procedure of DPZV. Based on (1), the objective is to collaboratively minimizing global objective function across all clients, each holds disjoint features of the same data records, and one server which holds the label data. The training procedure can be described in two iterative steps: (i) Client Update and Forward Communication: The sampled client computes local information and transmits it to the server, we define the client-server communication as forward communication. (ii) Server Update and Backward Communication: The server updates the global model and transmit global updates to the client, we define the server-client communication as backward communication.
@@ 1 Initialize $w_{0},w_{m}$ and set $t,t_{m}=0$ for all parties
3       Sample ready-to-update client $m\in\{1,\ldots,M\}$.
4       Client $m$ samples a mini-batch ${\mathcal{B}}_{m}^{t_{m}}$ with corresponding IDs ${\mathcal{I}}_{m}^{t_{m}}$.
5       Client $m$ computes perturbed local embeddings $\{h_{m,i}^{t_{m}+},h_{m,i}^{t_{m}-}\}_{i\in{\mathcal{I}}_{m}^{t_{m}}}$ according to (6).
6       (Forward) Client transmits embeddings to server.
7       Server computes $\Delta_{m}^{t}$ according to (7) and (8), and updates via ZO Optimization.
@@
9      (Backward) Client receives $\Delta_{m}^{t}$.
10
11      Client performs local update using (9).
12
13Client update local time stamp $t_{m}=t$.
14
@@
Training Procedure. Each client $m$ maintaining its own local communication round $t_{m}$, while the server tracks a global round $t$. Whenever the server receives information from a client, it increments $t$. Upon receiving an update from the server, the client synchronizes by setting $t_{m}=t$, capturing the latest state. This asynchronous mechanism allows clients to progress without waiting for stragglers, improving throughput and minimizing idle time. The server maintains a copy of the latest local embeddings $\tilde{h}^{t}_{m,i}$ for all clients $m\in[M]$ and data samples $\xi_{i}\in{\mathcal{D}}$. Due to the asynchronous nature of the algorithm, these copies may be stale, as they do not always reflect the most up-to-date model parameters of the clients. Let $\tilde{t}_{m,i}$ denote the client time when the server last updated $\tilde{h}^{t}_{m,i}$. The delay at server communication round $t$ can then be expressed as $\tau_{m}^{t}=t_{m}-\tilde{t}_{m,i},$ where the delayed model parameters are defined as:
$\textstyle\tilde{\chi}^{t}=\{w_{1}^{t_{1}-\tau_{1}^{t}},\dots,w_{M}^{t_{M}-\tau_{M}^{t}}\},\quad\tilde{{\bm{w}}}^{t}=\{w_{0}^{t},\tilde{\chi}^{t}\}.$ (5)
Local Embedding Perturbation. For each global iteration $t$, a client $m$ is activated, and it samples a mini-batch ${\mathcal{B}}_{m}^{t_{m}}\in{\mathcal{D}}$ and the corresponding IDs ${\mathcal{I}}_{m}^{t_{m}}$. To approximate gradients via zeroth-order finite differences, client $m$ computes two perturbed local embeddings for the mini-batch,
$\textstyle\{h_{m,i}^{t_{m}+}\}_{i\in{\mathcal{I}}_{m}^{t_{m}}}=\{h(w_{m}^{t_{m}}+\lambda_{m}{\bm{u}}_{m}^{t_{m}};\xi_{m,i})\}_{i\in{\mathcal{I}}_{m}^{t_{m}}},$
$\textstyle\{h_{m,i}^{t_{m}-}\}_{i\in{\mathcal{I}}_{m}^{t_{m}}}=\{h(w_{m}^{t_{m}}-\lambda_{m}{\bm{u}}_{m}^{t_{m}};\xi_{m,i})\}_{i\in{\mathcal{I}}_{m}^{t_{m}}}$ (6)
Server Side ZO Computation. After the server receives the local embeddings $h_{m,i}^{t_{m}}$ from client $m$, it updates its embeddings copy and do computation. For each embedding pair $\{h_{m,i}^{t_{m}+},h_{m,i}^{t_{m}-}\}$, it computes the difference in loss function ${\mathcal{L}}$ caused by perturbation, divided by the smooth parameter $\lambda_{m}$. Specifically, we define [^1] We slightly abuse the notation, and define $\tilde{f}(w_{0},h_{m,i}^{t_{m}\pm};y_{i})={\mathcal{L}}(w_{0},\tilde{h}^{t}_{1,i},\dots,h_{m,i}^{t_{m}\pm},\dots,\tilde{h}^{t}_{M,i};y_{i})$. where we treat $\tilde{f}$ as a function of the server parameter $w_{0}$ and the perturbed local embeddings.:
$\textstyle\delta_{m,i}^{t,t_{m}}=\frac{\tilde{f}(w_{0},h_{m,i}^{t_{m}+};y_{i})-\tilde{f}(w_{0},h_{m,i}^{t_{m}-};y_{i})}{\lambda_{m}},$ (7)
@@
$\textstyle\Delta_{m}^{t}=\frac{1}{B}\sum_{i\in{\mathcal{I}}_{m}^{t_{m}}}\mathrm{clip}_{C}(\delta_{m,i}^{t,t_{m}})+z_{m}^{t}.$ (8)
The server then performs a backward communication and sends $\Delta_{m}^{t}$ to client $m$. The server then updates its global model through two possible operations: $i)$ ZO Optimization, ${\bm{w}}_{0}\leftarrow{\bm{w}}_{0}-\eta_{0}g(w_{0};\xi_{i})$ (defined in (3)); or $ii)$ stochastic gradient descent (SGD), ${\bm{w}}_{0}\leftarrow{\bm{w}}_{0}-\eta_{0}\nabla_{w_{0}}F({\bm{w}};\xi_{i})$, depending on the constraints on computation resources. The adopted ZO update methodology is explained in Section 3. On the client side, upon receiving $\Delta_{m}^{t}$, client $m$ updates its local parameter ${\bm{w}}_{m}$ with learning rate $\eta_{m}$:
@@
5 Convergence Analysis
In this section, we provide the convergence analysis for DPZV. The detailed proofs can be found in Appendix B. For brevity, we define the following notations: $F^{t}=F({\bm{w}}^{t}):=F({\bm{w}}^{t};{\mathcal{D}},{\bm{Y}})$, and $f({\bm{w}};\xi_{i})$ as defined in (2). We make the following standard assumptions  [^2] We claim that 5.1 and 5.2 are standard in VFL and ZO literature[33, 3]. We follow [38] to make the $\ell$-Lipschitz assumption in order to bound the probability of clipping. 5.3 is common when dealing with asynchronous participation [4], and can be satisfied when the activations of clients follow independent Poisson processes.:
Assumption 5.1 (Properties of loss function).
@@ 5 Convergence Analysis
Assumption 5.2 (System boundedness).
@@
Assumption 5.3 (Independent Participation).
Under an asynchronous update system, the probability of one client participating in one communication round is independent of other clients and satisfies: (client $\mathbb{P}(\text{client }m\text{ uploading})=q_{m}.$
@@
Under 5.1-5.3. Define ${\mathcal{F}}=\mathbb{E}[F^{0}-F^{T}]$. If we denote $q_{*}=\min_{m}q_{m}$, $d_{*}=\max_{m}d_{m}$ where $d_{m}$ represent the dimension of model parameters on device $m$, and let all step sizes satisfy: $\eta_{0}=\eta_{m}=\eta\leq\min\{\frac{1}{\sqrt{Td_{*}}},\frac{B}{4L(B+8d_{0})+8\gamma_{1}(2d_{m}+B)}\}$, let the smoothing parameter $\lambda$ satisfy: $\lambda\leq\frac{1}{Ld\sqrt{T}}$, and let the clipping level $C$ satisfy: $C\geq\max\left\{0,\frac{1}{2}L\lambda d-\ell\sqrt{8\log(2\sqrt{2\pi})}\right\},$
then for any given global iteration $T\geq 1$, we have the following upper bound on the gradient norm:
$\textstyle\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[\mbox{$\left\lVert\nabla_{{\bm{w}}}F({\bm{w}}^{t})\right\rVert$}^{2}\right]\leq\mathcal{O}\bigg{(}\frac{{\mathcal{F}}\sqrt{d_{*}}}{\sqrt{T}}\textstyle+\frac{d_{*}}{T}+\frac{(\sigma_{s}^{2}/B+\sigma_{dp}^{2})\sqrt{d_{*}}}{\sqrt{T}}+\frac{C^{2}\sqrt{d_{*}}}{(\exp(C^{2})-1)B\sqrt{T}}\bigg{)},$ (10)

### v3

@@ III-B Differential Privacy and DP-SGD
One of the most widely adopted private training algorithms is Differentially Private Stochastic Gradient Descent (DP-SGD), which operates by clipping an intermediate result and injecting Gaussian noise during each model update. Specifically, given a minibatch $B_{t}$ at iteration $t$, DP-SGD performs the following steps:
  1. 1.
    Clip individual gradients:
    $\bar{g}_{i}=\nabla\ell(\theta_{t},x_{i}),\quad\tilde{g}_{i}=\bar{g}_{i}/\max\left(1,\frac{\|\bar{g}_{i}\|_{2}}{C}\right)$
  2. 2.
    Add Gaussian noise:
    $\tilde{g}_{t}=\frac{1}{|B_{t}|}\left(\sum_{i\in B_{t}}\tilde{g}_{i}+\mathcal{N}(0,\sigma^{2}C^{2}\mathbf{I})\right)$
  3. 3.
    Update model: $\theta_{t+1}=\theta_{t}-\eta_{t}\tilde{g}_{t}$
III-C Challenges and Motivation
@@
Existing approaches [chen2020vafl, xie2024improving] integrate DP with first order VFL by injecting DP noise in the forward embedding. In centralized or horizontally federated learning, DP-SGD achieves privacy by adding noise to the backward gradients. However, extending these strategies to ZO optimization introduces new challenges. First, due to the distributed nature of VFL, DP guarantees must be enforced during the downlink process between the cloud and devices, so as to prevent malicious devices from mounting label inference attacks. In both cases, the noise must match the dimensionality of the protected vector (embedding or gradient), often resulting in high-dimensional noise. This situation is further exacerbated by the second factor: in ZO methods, gradients are approximated through multiple forward evaluations, making them inherently noisy. Adding a high-dimensional noise vector as in standard DP-SGD further amplifies this inaccuracy.
@@
A key advantage of adopting ZO optimization, however, lies in its communication efficiency in the backward pass: only a scalar value, rather than a full gradient vector, needs to be exchanged. This observation enables a more efficient privacy mechanism—injecting scalar noise rather than high-dimensional noise. Building on this insight, we propose Differentially Private Zeroth-Order Vertical Federated Learning (DPZV), a framework that achieves tunable differential privacy through calibrated scalar noise injection. This design allows for a favorable privacy-utility trade-off in ZO-based VFL, particularly under tight privacy budgets. Unlike first-order VFL methods, which typically add noise to forward embeddings, DPZV introduces noise in the backward pass, targeting the most exploited leakage pathway in VFL: label inference from gradients. A broader discussion of privacy implications is provided in Section VI.
IV Methodology
@@
In this section, we describe the overall training procedure of DPZV. Based on (1), the objective is to collaboratively minimize the global objective function across all devices. The devices hold disjoint features of the same data records, and the server holds the label data. The training procedure proceeds in two iterative steps.
IV-A Device Update and Forward Communication
The sampled devices compute local information and transmit it to the server. We define the device-server communication as forward communication.
Training Procedure. Each device $m$ maintaining its own local communication round $t_{m}$, while the server tracks a global round $t$. Whenever the server receives information from a device, it increments $t$. Upon receiving an update from the server, the device synchronizes by setting $t_{m}=t$, capturing the latest state. This asynchronous mechanism allows devices to progress without waiting for stragglers, improving throughput and minimizing idle time. The server maintains a copy of the latest local embeddings $\tilde{h}^{t}_{m,i}$ for all devices $m\in[M]$ and data samples $\xi_{i}\in{\mathcal{D}}$. Due to the asynchronous nature of the algorithm, these copies may be stale, as they do not always reflect the most up-to-date model parameters of the devices. Let $\tilde{t}_{m,i}$ denote the device time when the server last updated $\tilde{h}^{t}_{m,i}$. The delay at server communication round $t$ can then be expressed as $\tau_{m}^{t}=t_{m}-\tilde{t}_{m,i},$ where the delayed model parameters are defined as:
$\tilde{\chi}^{t}=\{w_{1}^{t_{1}-\tau_{1}^{t}},\dots,w_{M}^{t_{M}-\tau_{M}^{t}}\},\quad\tilde{{\bm{w}}}^{t}=\{w_{0}^{t},\tilde{\chi}^{t}\}.$ (5)
Local Embedding Finite Differences. For each global iteration $t$, a device $m$ is activated, and it samples a mini-batch ${\mathcal{B}}_{m}^{t_{m}}\in{\mathcal{D}}$ and the corresponding IDs ${\mathcal{I}}_{m}^{t_{m}}$. To approximate gradients via zeroth-order finite differences, device $m$ computes two perturbed local embeddings for the mini-batch,
$\{h_{m,i}^{t_{m}+}\}_{i\in{\mathcal{I}}_{m}^{t_{m}}}=\{h(w_{m}^{t_{m}}+\lambda_{m}{\bm{u}}_{m}^{t_{m}};\xi_{m,i})\}_{i\in{\mathcal{I}}_{m}^{t_{m}}},$
$\{h_{m,i}^{t_{m}-}\}_{i\in{\mathcal{I}}_{m}^{t_{m}}}=\{h(w_{m}^{t_{m}}-\lambda_{m}{\bm{u}}_{m}^{t_{m}};\xi_{m,i})\}_{i\in{\mathcal{I}}_{m}^{t_{m}}}$ (6)
@@ 1 Initialize $w_{0},w_{m}$ and set $t,t_{m}=0$ for all parties
3   Sample ready-to-update device $m\in\{1,\ldots,M\}$.
4   device $m$ samples a mini-batch ${\mathcal{B}}_{m}^{t_{m}}$ with corresponding IDs ${\mathcal{I}}_{m}^{t_{m}}$.
5   device $m$ computes perturbed local embeddings:
$\{h_{m,i}^{t_{m}+}\}_{i\in{\mathcal{I}}_{m}^{t_{m}}}=\{h(w_{m}^{t_{m}}+\lambda_{m}{\bm{u}}_{m}^{t_{m}};\xi_{m,i})\}_{i\in{\mathcal{I}}_{m}^{t_{m}}},$
$\{h_{m,i}^{t_{m}-}\}_{i\in{\mathcal{I}}_{m}^{t_{m}}}=\{h(w_{m}^{t_{m}}-\lambda_{m}{\bm{u}}_{m}^{t_{m}};\xi_{m,i})\}_{i\in{\mathcal{I}}_{m}^{t_{m}}}$
(Forward) Device transmits embeddings to server.
6   Server computes $\Delta_{m}^{t}$, and updates via ZO Optimization: $\delta_{m,i}^{t,t_{m}}=\frac{\tilde{f}(w_{0},h_{m,i}^{t_{m}+};y_{i})-\tilde{f}(w_{0},h_{m,i}^{t_{m}-};y_{i})}{\lambda_{m}},$ $\Delta_{m}^{t}=\frac{1}{B}\sum_{i\in{\mathcal{I}}_{m}^{t_{m}}}\mathrm{clip}_{C}(\delta_{m,i}^{t,t_{m}})+z_{m}^{t}.$
7 (Backward) Device receives $\Delta_{m}^{t}$.
@@
9  Device performs local update:
${\bm{w}}_{m}\leftarrow{\bm{w}}_{m}-\eta_{m}\Delta_{m}^{t}{\bm{u}}_{m}^{t_{m}}.$
10Device update local time stamp $t_{m}=t$.
@@ 9  Device performs local update:
IV-B Server Update and Backward Communication
The server updates the global model and transmit global updates to the device. We define the server-device communication as backward communication.
Server Side ZO Computation. After the server receives the local embeddings $h_{m,i}^{t_{m}}$ from device $m$, it updates its embeddings copy and do computation. For each embedding pair $\{h_{m,i}^{t_{m}+},h_{m,i}^{t_{m}-}\}$, it computes the difference in loss function $\mathcal{L}$} caused by perturbation, divided by the smooth parameter $\lambda_{m}$. Specifically, we define [^1] We slightly abuse the notation, and define $\tilde{f}(w_{0},h_{m,i}^{t_{m}\pm};y_{i})={\mathcal{L}}(w_{0},\tilde{h}^{t}_{1,i},\dots,h_{m,i}^{t_{m}\pm},\dots,\tilde{h}^{t}_{M,i};y_{i})$. where we treat $\tilde{f}$ as a function of the server parameter $w_{0}$ and the perturbed local embeddings.:
$\delta_{m,i}^{t,t_{m}}=\frac{\tilde{f}(w_{0},h_{m,i}^{t_{m}+};y_{i})-\tilde{f}(w_{0},h_{m,i}^{t_{m}-};y_{i})}{\lambda_{m}},$ (7)
@@ IV-B Server Update and Backward Communication
$\Delta_{m}^{t}=\frac{1}{B}\sum_{i\in{\mathcal{I}}_{m}^{t_{m}}}\mathrm{clip}_{C}(\delta_{m,i}^{t,t_{m}})+z_{m}^{t}.$ (8)
The server then performs a backward communication and sends $\Delta_{m}^{t}$ to device $m$. The server then updates its global model through two possible operations: $i)$ ZO Optimization, ${\bm{w}}_{0}\leftarrow{\bm{w}}_{0}-\eta_{0}g(w_{0};\xi_{i})$ (defined in (3)); or ii)ii) stochastic gradient descent (SGD), ${\bm{w}}_{0}\leftarrow{\bm{w}}_{0}-\eta_{0}\nabla_{w_{0}}F({\bm{w}};\xi_{i})$,
depending on the constraints on computation resources. The adopted ZO update methodology is explained in Sec III.
On the device side, upon receiving $\Delta_{m}^{t}$, device $m$ updates its local parameter ${\bm{w}}_{m}$ with learning rate $\eta_{m}$:
@@
V Convergence Analysis
In this section, we provide the convergence analysis for DPZV. For brevity, we define the following notations: $F^{t}=F({\bm{w}}^{t}):=F({\bm{w}}^{t};{\mathcal{D}},{\bm{Y}})$, and $f({\bm{w}};\xi_{i})$ as defined in (2). We make the following standard assumptions  [^2] Assumption V.1 and V.2 are standard in VFL and ZO literature [wang2024unified, castiglia2023flexible]. We follow [zhang2024dpzero] to make the $\ell$-Lipschitz assumption in order to bound the probability of clipping. Assumption V.3 is common when dealing with asynchronous participation [chen2020vafl], and can be satisfied when the activations of devices follow independent Poisson processes.:
Assumption V.1 (Properties of loss function).
@@ V Convergence Analysis
Assumption V.2 (System boundedness).
@@
Assumption V.3 (Independent Participation).
Under an asynchronous update system, the probability of one device participating in one communication round is independent of other devices and satisfies: $\mathbb{P}(\text{device }m\text{ uploading})=q_{m}.$
V-A Convergence Guarantee and Discussion
@@
Under assumption V.1-V.3, define ${\mathcal{F}}=\mathbb{E}[F^{0}-F^{T}]$. Denote $q_{*}=\min_{m}q_{m}$, $d_{*}=\max_{m}d_{m}$ where $d_{m}$ represent the dimension of model parameters on device $m$, let all step sizes satisfy: $\eta_{0}=\eta_{m}=\eta\leq\min\{\frac{1}{\sqrt{Td_{*}}},\frac{B}{4L(B+8d_{0})+8\gamma_{1}(2d_{m}+B)}\}$, let the smoothing parameter $\lambda$ satisfy: $\lambda\leq\frac{1}{Ld\sqrt{T}}$, and let the clipping level $C$ satisfy: $C\geq\max\{0,\frac{1}{2}L\lambda d-\ell\sqrt{8\log(2\sqrt{2\pi})}\}.$ Then, for any given global iteration $T\geq 1$, we have the following upper bound on the gradient norm:
$\leq$ $\mathcal{O}\bigg(\frac{{\mathcal{F}}\sqrt{d_{*}}}{\sqrt{T}}+\frac{d_{*}}{T}+\frac{(\sigma_{s}^{2}/B+\sigma_{dp}^{2})\sqrt{d_{*}}}{\sqrt{T}}$
$+\frac{C^{2}\sqrt{d_{*}}}{(\exp(C^{2})-1)B\sqrt{T}}\bigg),$ (10)